OpenStudy (waheguru):

−12 = 3 − 2k − 3k Can someone show how to solve this step by step

mathslover (mathslover):

u want to solve for k right ?

OpenStudy (waheguru):

−12 = 3 − 2k − 3k -12=2-5k now what do i do ?

OpenStudy (waheguru):

I ment -12=3-5k

OpenStudy (anonymous):

-12=3-2k-3k -12=3-5k -15=-5k 15=5k 3=k

mathslover (mathslover):

\[\huge{-12=2-5k}\] \[\huge{-12-2=-5k}\] \[\huge{-15=-5k}\] \[\huge{5k=15}\] \[\huge{k=\frac{15}{5}=3}\]

OpenStudy (waheguru):

@ivan514 why do you subtract 12-3 rather than add

OpenStudy (waheguru):

−12−3=−5k At this step why do we subtract ?

OpenStudy (anonymous):

-12=3-5k isolate variable therefore (-3)-12=3-5k(-3) that will give you isolated variable on right side.

OpenStudy (waheguru):

So when we move a number we give it the opposite operation right?

mathslover (mathslover):

yes .. transposing

OpenStudy (waheguru):

transposing ?

mathslover (mathslover):

\[\huge{x+a=b}\] \[\huge{x+a-a=b-a}\] subtracting both sides by a \[\huge{x=b-a}\] hence a gets opposite sign there in right hand side transposing is the name of this process

OpenStudy (waheguru):

Thanks

mathslover (mathslover):

ur welcome and best of luck

OpenStudy (waheguru):

=)